Dinitrogen tetroxide decomposes to nitrogen dioxide: N2O4(g)→2NO2(g) ΔrH∘=55.3kJmol−1 At 298 K, a reaction vessel initially contains 0.100 bar of N2O4. When equilibrium is reached, 58% of the N2O4 has decomposed to NO2.

What percentage of N2O4 decomposes at 389 K? Assume that the initial pressure of N2O4 is the same (0.100 bar).

Please use bar and not atm.
Answer #1

We are given that are given that initial po essure of Noy = 0.100 bar At equillibrium N, Ou decomposes 58%.. Firetty we will

Given that, Mn = 55.34 Ime at T = 298k This value is for prenure constant K . Now we cwill colulate pressure constant kg, at

)! k DR (No = (a (0.1 -x) NOY (93) = 59.17 59.17 (0.1 -0 40 + 59.17 - 5.917 = 0 Solving for u, where a=4, b= 59.17., c= -5.9